OpenStudy (anonymous):

Can somebody help me with this? 2x + 3y = 3 and 3x – 2y = 11 Possible answers: (1, –4) (3, –1) (–1, –7) (–3, 3) I choose (3,-1) but not sure.. Thank you for your help.

OpenStudy (anonymous):

You are absolutely right.

OpenStudy (anonymous):

2X+3y=3 3X-2Y=11 Using elimination method 3(2x+3y=3) 6x+9y=9 2(3x-2y=11) 6x-4y=22 subtracting the second equation from the first one 13y=-13 therefore y=-1 substituting -1 in equation two 3x-2(-1)=11 3x+2=11 3x= 9 x=3 therefore -1 and 3 are the correct values of y and x respectively.